# Amputation-imputation based generation of synthetic tabular data for ratemaking

### v1

@@ 4.2 Tested methods and formulas
  • CTGAN (CTGAN) – implementation of Xu et al., (2019) available in the sdv library for Python;
@@
  • Wasserstein GAN with gradient penalty (COTE_MC_WGAN_GP) – implementation of the method graciously provided by Cote et al., (2020);
  • GAN-based tabulator (JCS_TABULATOR) – tabulator as proposed and implemented by Neves et al., (2022).
@@
The second group of synthetic-data generators we test are the imputation-amputation methods based on MICE. One of the considered approaches is tabulator proposed Neves et al., (2022), but we use MICE as the imputation engine instead of GAIN-based models suggested there. In our “MICE-tabulator”, we have $5$ sequential amputation-imputation iterations, each of each amputates randomly selected $20\%$ of cells and imputes them to generate synthetic data points. The other two methods have a different amputation logic. First, we create a copy of the training data, where we remove $75\%$ of cell values. Second, we append this data with missing values to the original training data and apply MICE to impute the missing values. The data where missing cells were present, is treated as the generated data. In this way, we end up with generated data where 25% of the data comes from the original training dataset and 75% is synthetic. Hereinafter, we refer to this approach as the “MICE method”. The third approach extends the MICE method by the second iteration, in which the cells that were not amputated-imputed in the first iteration are amputated and imputed via MICE. This leads to a data set that is fully synthetic. In summary, the MICE-based methods we consider are
@@
  • MICE method (MICE) – the method described above and implemented using the R-package mice;
@@
  • MICE with all synthetic data (MICE_ALL_SYN) – the modification of the MICE method, where in an additional step the remaining $25\%$ of the original data is imputed via MICE;
@@
  • MICE tabulator (MICE_TABULATOR) – the tabulator setup from Neves et al., (2022) with $5$ iterations and MICE with RFs as the imputation engine.
@@
After testing the above-described groups of methods, we noticed that GAN-based approaches learned the univariate distribution of numeric variables worse than MICE-based methods and had challenges with high-cardinality categorical variables. Due to this finding, we decided to consider in our case studies two hybrid models. One of them combines a CTGAN and a MICE component, namely a CTGAN first generates synthetic data and then $75\%$ of synthetic numeric variable values and imputed amputated and imputed using MICE. Note that also for this CTGAN-MICE model, the data with removed cell values was combined with the original training data to impute the missing numerical column values. The second hybrid model that we tested enhances CTGAN-MICE by using AEs to encode categorical variables before applying CTGAN-MICE and then using clustering with AEs to decode synthetic AE-encoded categorical variables. So, the hybrid approaches we study are:
@@
Lastly, we also test the method proposed in Jamotton and Hainaut, (2024), where variational autoencoders (VAEs) are used for data generation. We use the implementation graciously shared with us by the authors of Jamotton and Hainaut, (2024). The monospaced shorthand for this method is VAE_JAMOTTON.
As described in Subsection 4.1, artificial claim counts are generated for this case study such a way that we know the true relationship between the response variable and the descriptive variables. The following formula for a linear relationship is used:
@@
The claim counts simulated using (4.3) and (4.2) lead to the average claim frequency of $5.37\%$ and $6.04\%$ respectively. These numbers are close to the average claim frequencies that we have observed in the industry when modeling MTPL claims.
@@ 4.3 Case study results
With our case study, we evaluated a total of $10$ different approaches to synthetic data generation. However, we display the results only for $9$ of them, as the tested tabulator by Neves et al., (2022) did not generate any number of claims, and, therefore, it could not be fully evaluated.
@@
Data-specific metrics are MAE/MAPE for numeric, categorical and pair-wise combinations of columns and MAE/MAPE for correlation structure within numeric columns of the generated data. Lastly, our subjective rank of the method’s ease of use is also provided.
Note that every synthetic dataset is uniquely generated 5 times ($n_{E}=5$) and all of the following results are presented as simple averages across all generated datasets for a given method (except for $M_{1}$ and $M_{2}$ metrics, as these are already averaged across $n_{E}$). For data-specific metrics, every column is taken as equally important, so the shown metric is an average taken over all columns or column combinations.
The final results for the evaluation of methods are presented in Tables 1 and 2. Each cell shows the rank of the method with respect to the metric shown in the column. Additionally, the table is color matched based on the rank of the method, and in the case of ties, with respect to metric, all cells are given the same highest rank available. Note that in all cases, the ”training” method is shown as the reference value for different metrics.
@@
(6,6)
4
(1,2)
5
(0,32545)
5
(0,23829)
9
(0,00303)
3
(0,40646)
10
(0,03651)
9
(1,07215)
7
(0,00429)
7
(0,0809)
9
(0,02964)
3
(0,32234)
9
COTE_MC_WGAN_GP
8
(155,80504)
9
(152,74518)
1
8
7
(0,33124)
7
(0,23896)
10
(0,00412)
10
(82,61544)
6
(0,01052)
6
(0,48478)
8
(0,00444)
8
(0,08287)
10
(0,04028)
10
(445,05061)
10
CTGAN_MICE
10
(393,44698)
10
(383,62816)
1
6
(1,4)
9
(0,34255)
9
(0,23977)
8
(0,00292)
8
(8,9032)
2
(0,00543)
5
(0,44177)
9
(0,00678)
9
(0,11463)
7
(0,0147)
8
(42,48769)
6
CTGAN_WITH_AE_MICE
9
(143,03862)
8
(153,73562)
6
(6,8)
6
(1,4)
8
(0,33266)
8
(0,23937)
5
(0,00199)
7
(8,25317)
3
(0,00569)
8
(0,71934)
3
(0,0016)
5
(0,05566)
6
(0,00974)
7
(38,72062)
8
CTGAN_WITH_AE
7
(125,10637)
6
(136,08475)
6
(6,8)
@@
Looking at Table 1, we can see that in no case does any method outperform training data. The differences between training data and synthetic data can be quite large for example for metric $M_{1}$ or $M_{2}$. However, when it comes to the overall accuracy of the model, the difference in coefficients does not result in a very significant difference in the accuracy of the predictions.
@@
We can see that the best-performing synthetic data generation methods are MICE and CTGAN. They show almost the same accuracy in terms of goodness-of-fit metrics, but MICE is better when it comes to the accuracy of model coefficients. The worst performing methods are CTGAN_MICE, CTGAN_WITH_AE and CTGAN_WITH_AE_MICE. All of these methods not only show worse accuracy in model coefficients but also the model predictions are worse compared to other methods.
@@
Almost all methods show consistency in the correct terms using the variable selection method. However, for all methods the variable selection included additional variables not present in the true model. Here, the worst performing method was MICE_TABULATOR, where automatic variable selection on average failed to include 1.4 correct variables and included, on average, 2 additional incorrect variables (the worst possible result) to the model. The best variable selection results were achieved by models trained on data using the CTGAN method, always selecting the correct variables and including only 1 additional incorrect variable on average.
@@
Similar observations can be seen when it comes to data-specific metrics. Here again, MICE performs the best across all data metrics. From Table 1, we conclude that different methods have different niches. For example, it seems MICE_TABULATOR is much better at categorical column generation, followed closely by CTGAN_WITH_AE and CTGAN_WITH_AE_MICE (since they share categorical variables). This possibly shows that the usage of autoencoders improves the generation of categorical columns. Additionally, we can observe that the addition of MICE to CTGAN-based methods might not be very consistent. MICE helps in terms of numerical column accuracy for CTGAN_WITH_AE_MICE, but makes it worse for CTGAN method. WGAN by Cote et al., (2020) shows overall the worst performance when it comes to data-specific metrics.
As for to the ease of use, MICE is one of the easiest methods to apply, since the implementation in the R package mice is straightforward and requires no additional implementation or data preparation besides the choice of parameters for the function. CTGAN-based methods seem to require more tinkering and also a change of environment to Python since the implementation of the CTGAN interface in R (through the package ctgan) in our experience is neither robust nor intuitive. Addition of autoencoders, although simple in idea, can be quite hard to implement since this requires additional steps with respect to data preparation, setting up and fitting the additional neural networks and finally clustering the generated vectors to get back to the correct categorical labels. The most complex implementations were those provided by Jamotton and Hainaut, (2024) and Cote et al., (2020) since they required the most additional work to make it work properly in our case study.
@@
(1,6)
6
(0,4)
3
(0,34166)
3
(0,25323)
3
(0,00073)
4
(0,45389)
7
(0,01121)
3
(0,37872)
9
(0,$01$)
9
(0,09023)
4
(0,00175)
5
(0,41926)
3
MICE_ALL_SYNTHETIC
8
(55,7844)
8
(49,90837)
7
(9,4)
7
10
(1,2)
5
(0,34909)
5
(0,25434)
4
(0,00111)
5
(0,4616)
5
(0,00967)
5
(0,38692)
4
(0,00242)
4
(0,02565)
2
(0,00102)
4
(0,3482)
3
MICE_TABULATOR
7
(53,21122)
6
(49,87283)
8
(8,4)
@@
From model-specific results, we see that coefficient distance metrics $M_{1}$ and $M_{2}$ have much smaller values compared to the linear case. The best methods based on coefficient difference metrics are again MICE and CTGAN, and the worst method seems to be COTE_MC_WGAN_GP and CTGAN_MICE. Notably, CTGAN_WITH_AE and CTGAN_WITH_AE_MICE show a big uplift in rankings while additional MICE-based methods lose ranks compared to the linear case. This possibly shows that MICE is less suited for data structures with inherent dependence, but overall performance still seems to be the best across the tested approaches.
Model goodness-of-fit metrics reflect very similar patterns as model coefficient metrics. However, MICE_ALL_SYNTHETIC shows a higher rank with respect to the goodness-of-fit metrics. The Poission deviance for it is notably smaller (0.00554) than that for MICE_TABULATOR.
The variable selection shows, however, much more varied results. Again, MICE and CTGAN methods show best results, where variable selection, on average, selected 15.4 and 15.6 correct terms. Beyond these two methods, the variable selection in the next best method selected, on average, 12 correct terms. Variable selection for VAE_JAMOTTON got the worst result, on average, selecting only 7.8 correct terms to the model. We can also see that for almost all methods, variable selection added additional incorrect terms, where only VAE_JAMOTTON picked on average 0.4 additional incorrect terms. These combined facts indicate that using VAE_JAMOTTON results in a small/lean model, which is good for the linear case, but worse for the interaction case.
From the data-specific results, we can observe almost the same rankings as we did for the linear case. This is expected since we did not introduce any additional inter-variable relationships besides the interaction relationship between response and the two covariates. The biggest difference is the fact that autoencoders did not improve the MAE/MAPE of categorical columns as significantly for CTGAN method as they had done in the linear case study. MICE is again clearly the leader when it comes to data-specific metrics, achieving top ranks in all metrics. However, analyzing synthetic data on a more granular level, we noticed that MICE-based methods can generate numeric values that ignore business constraints, e.g., generating observations with a driver age below $18$. Imposing such constraints may be investigated in a follow-up paper.
@@
Having discussed data generation fidelity, we now focus on data augmentation. For this, we look at $M_{1}$ metric depending on the proportion of synthetic data added. We observed that as we increase the proportion of additional synthetic data compared to the original training data, the value for $M_{1}$ metric increases. The only case where we could observe a decrease in a metric was for coefficient distance metrics $M_{1}$ for the method MICE with an additional synthetic data proportion of $0.2$. In this case, the augmented data achieved the $M_{1}$ metric value of $0.78717$ while the corresponding training data metric was $0.86047$. Since this is the only result where this happened, it is not enough to say that augmentation of the original freMTPL2freqdata might help when it comes to better accuracy of coefficients.
The general relationship between the increase of additional synthetic data proportion and metric value can be seen in Figure 2. The shown figure is for metric $M_{1}$. As the synthetic data proportion increases, we can see a clear increasing trend for $M_{1}$ metric value for the given method. Based on the way $M_{1}$ metric is calculated and how GLM relies on the underlying data to compute the coefficient values, we estimate that if we keep on increasing the proportion of additional synthetic data, then we will asymptotically reach the metric values which we can compute for case where we use only synthetic data. These pure synthetic data results are shown on the right side of the plot.

### v2

@@ 4.2 Tested methods and formulas
  • CTGAN (CTGAN) – implementation of Xu et al. (2019) available in the sdv library for Python;
@@
  • Wasserstein GAN with gradient penalty (COTE_MC_WGAN_GP) – implementation of the method graciously provided by Côté et al. (2025);
  • GAN-based tabulator (JCS_TABULATOR) – tabulator as proposed and implemented by Neves et al. (2022).
@@
The second group of synthetic-data generators we test are the imputation-amputation methods based on MICE. One of the considered approaches is the tabulator proposed by Neves et al. (2022), but we use MICE with RFs as the imputation engine instead of the GAIN-based models suggested there. In our MICE-based tabulator, we have $5$ sequential amputation-imputation iterations, each of which sets random $20\%$ of cells to missing and imputes them using MICE-RF to generate synthetic data points. The other two methods have a different amputation logic. First, we create a copy of the training data, where we set $75\%$ of cell values to missing. Second, we append this data with missing values to the original training data and apply MICE-RF to impute the missing values. The data with missing cells is treated as generated data. In this way, we end up with generated data where 25% of the data comes from the original training dataset, and 75% is synthetic. Although the resulting dataset is partially synthetic, the probability of disclosing confidential identifiable information or of having the same row in both the synthetic data and the original data should be negligible due to the completely-at-random way of creating missingness in the original data. Hereinafter, we refer to this approach as the “MICE partially synthetic method”. The third approach extends the MICE partially synthetic method by the second iteration, in which the cells that were not imputed in the first iteration (the remaining quarter of the original data) are set to missing and then imputed via MICE. This results in a fully synthetic dataset. This method is referred to as “MICE fully synthetic method”. For each of the approaches described in this paragraph, we use only one synthetic dataset for evaluating data fidelity and data utility. This is different from the original idea of MICE for statistical inference (see Subsection 3.1, Figure 1), where multiple different imputed datasets are created and then combined to study a statistical object with the help of Rubin’s rules. In summary, the MICE-based methods we explore and benchmark in our case study are:
@@
  • MICE partially synthetic method (MICE_PART_SYN) – the method described above and implemented using the R-package mice;
@@
  • MICE fully synthetic method (MICE_FULL_SYN) – the modification of the partially synthetic method, where in an additional step the remaining $25\%$ of the original data is imputed via MICE;
@@
  • MICE tabulator (MICE_TABULATOR) – the tabulator setup from Neves et al. (2022) with $5$ iterations and MICE with RFs as the imputation engine.
  • MICE as per Volker-Vink (MICE_VV) – amputation-imputation procedure utilizing MICE columnwise as described in Volker and Vink (2021).
Before we move on to other synthetic data generators in our case study, we clarify the different types of “iterations” in MICE-based methods. There are amputation-imputation iterations and there are MICE iterations. The former iteration type means that, first, a not yet imputed part of the training data is set to missing and, second, the respective missing values are imputed (using MICE-RF in our case study). A MICE iteration should be understood as a sequence of actions in Steps 2-5 in Figure 1 done one single time, i.e., once for each variable creating missingness, training an RF that uses other variables to predict the variable with missingness, and replacing introduced missing values by RF predictions. For example, in our case study, the MICE-tabulator approach has $5$ amputation-imputation iterations, the MICE partially synthetic method has $1$ amputation-imputation step, and the MICE fully synthetic method has $2$ amputation-imputation steps. In the MICE-based generation as per Volker and Vink (2021), there is no explicit creation of missingness before training imputation models. Instead, for each column, an imputation model is trained without randomly deleting observed values. Then each observed value is set to missing and then replaced by the respective prediction from the trained imputation model. The resulting imputed column is then used as the predictor for further imputation in the chain of columns. Since a single pass over the original dataset is performed only once to obtain a fully synthetic dataset, there are in total $n_{C}$ amputation-imputation iterations.
@@
After testing the above-described groups of methods, we noticed that GAN-based approaches learned the univariate distributions of numeric variables worse than MICE-based methods and had challenges with high-cardinality categorical variables. Due to this finding, we decided to consider two hybrid models. One of them combines a CTGAN and a MICE component, namely a CTGAN first generates synthetic data and then $75\%$ of the synthetic numeric variable values are set to missing and then imputed using MICE. Note that also for this CTGAN-MICE model, the data with removed cell values was combined with the original training data to impute the missing numerical column values. The second hybrid model that we tested enhances CTGAN-MICE by using AEs to encode categorical variables before applying CTGAN-MICE and then using clustering with AEs to decode synthetic AE-encoded categorical variables. So, the hybrid approaches we study are:
@@
Lastly, we test the method proposed in Jamotton and Hainaut (2024), which uses variational autoencoders (VAEs) for data generation. We use the implementation graciously shared with us by the authors of Jamotton and Hainaut (2024). The monospaced shorthand for this method is VAE_JAMOTTON.
As described in Subsection 4.1, artificial claim counts are generated for this case study in such a way that we know the true relationship between the response variable and the descriptive variables. The following formula for a linear relationship is used:
@@
The claim counts simulated using (4.3) and (4.2) lead to the average claim frequency of $5.37\%$ and $6.04\%$, respectively. These numbers are close to the values observed in practice.
We acknowledge that the chosen linear and interaction formulas follow the setup of generalized linear models, which is the core basis of insurance ratemaking in practice. The topic of more complex dependence structures is left as an avenue for further research.
@@ 4.3 Case study results
With our case study, we evaluated a total of $11$ different approaches to synthetic data generation. However, we display the results only for $10$ of them, as the tested tabulator by Neves et al. (2022) did not generate any number of claims, and, therefore, it could not be fully evaluated.
@@
Data-specific metrics are MAE/MAPE for numeric, categorical and pair-wise combinations of columns and MAE/MAPE for correlation structure within numeric columns of the generated data. Lastly, our subjective ranking of the method’s ease of use is also provided.
Note that every synthetic dataset is uniquely generated 5 times ($n_{E}=5$) and all of the following results are presented as simple averages across all generated datasets for a given method (except for $M_{1}$ and $M_{2}$ metrics, since these incorporate $n_{E}$). For data-specific metrics, each column is taken as equally important, so the metrics reflect the average of all columns or combinations.
The final results for the evaluation of methods are presented in Tables 1 and 2. Each cell shows the rank of the method with respect to the metric shown in the column. Additionally, the table is color-matched based on the rank of the method, and in the case of ties, with respect to the metric, all cells are given the same highest rank available. Note that in all cases, the ”training” method is shown as the reference value for different metrics.
Before looking at the results, it is important to comment that in our case study all generators were trained using the University of Tartu high-performance computation cluster (University of Tartu (2018)). Because this cluster does not inherently have hardware stability, we were unable to directly compare fitting and generation times. However, we observed that MICE-based and CTGAN-based methods usually took the least time (around 3 hours for model fitting and data generation), while more complex custom implementations, such as COTE_MC_WGAN_GP, took significantly longer for model fitting and generation (about 10 hours in total).
@@
Looking at Table 1, we can see that in no case does any method outperform training data. The differences between training data and synthetic data can be quite large for example for metric $M_{1}$ or $M_{2}$. However, when it comes to overall model accuracy, the difference in coefficients does not translate into a significant difference in prediction accuracy.
@@
We can see that the best-performing synthetic data generation methods are MICE_PART_SYN and MICE_FULL_SYN. They show almost the same accuracy in terms of goodness-of-fit metrics, but MICE_PART_SYN is better when it comes to the accuracy of model coefficients. The worst performing methods are CTGAN, CTGAN_WITH_AE_MICE, COTE_MC_WGAN_GP and CTGAN_WITH_AE. These methods have higher model coefficient metric values and worse accuracy in model predictions.
@@
Almost all methods show consistency in the correct terms using the variable selection method. However, for all methods, the variable selection included additional variables not present in the true model. Here, the worst-performing method was CTGAN_MICE, where automatic variable selection failed to include 1.4 correct variables, on average, and included, on average, 2 additional incorrect variables (the worst possible result). The best variable selection results were achieved by models trained on data using the MICE_PART_SYN and MICE_FULL_SYN methods, which always selected the correct variables and included only 1 additional incorrect variable on average. Notably, MICE_VV included, on average, only $0.4$ incorrect variables.
@@
Similar observations can be seen for data-specific metrics. Here again, MICE_PART_SYN overall performs the best across all data metrics. From Table 1, we conclude that different methods have different niches. For example, it seems MICE_TABULATOR is much better at categorical column generation, followed closely by CTGAN_WITH_AE and CTGAN_WITH_AE_MICE (since they share categorical variables). This suggests that the use of autoencoders improves the generation of categorical columns. Additionally, we observe that adding MICE to CTGAN-based methods slightly improves the generation of numeric columns, as both CTGAN_MICE and CTGAN_WITH_AE_MICE perform better than their non-augmented counterparts. WGAN by Côté et al. (2025) shows the worst overall performance on data-specific metrics.
We also qualitatively assess the usability of each tested generator based on three criteria: intuitiveness (conceptual clarity), operational complexity (training and usage pipelines), and deployment overhead (environment setup, debugging, and code adaptation).
We find MICE-based methods to be the most accessible, primarily due to their streamlined implementation in the R package mice, which requires minimal setup beyond standard installation. Notably, mice outputs by default synthetic data without exporting the trained imputation models.
In contrast, CTGAN-based methods require additional setup time and more extensive data preprocessing (e.g., matrix fragmentation, type casting). For practical reasons, these experiments were conducted in Python, as this environment was more compatible with our computational workflow than the available R implementation.
Incorporating AEs, while conceptually straightforward, increases implementation complexity due to the need for supplementary data preparation, neural network specification, and post-hoc clustering to recover categorical labels.
Finally, the custom implementations kindly provided by the authors of Jamotton and Hainaut (2024) and Côté et al. (2025) represent highly specialized, advanced architectures. Consequently, integrating these custom-built models into a standardized, comparative computational framework naturally necessitates a higher degree of code adaptation and environment configuration.
@@
CTGAN_MICE
@@
From the model-specific results, we see that the coefficient distance metrics $M_{1}$ and $M_{2}$ have much smaller values than in the linear case. The best methods based on coefficient difference metrics are again MICE_PART_SYN and MICE_FULL_SYN, and the worst method seems to be COTE_MC_WGAN_GP and CTGAN. Notably, CTGAN_WITH_AE and CTGAN_WITH_AE_MICE show a big uplift in rankings while additional MICE-based methods lose ranks compared to the linear case. This suggests that MICE is less suited to data structures with inherent dependence, but overall performance still seems best across the tested approaches.
Model goodness-of-fit metrics yield slightly different results than model coefficient metrics. MICE-based methods lead in terms of these metrics, while GAN and VAE-based methods show worse results. Notably, CTGAN_WITH_AE shows much better results when compared to other GAN-based generation methods.
The variable selection, however, shows much more varied results. Again, MICE_PART_SYN and MICE_FULL_SYN methods show the best results, where variable selection, on average, selected 15.4 and 15.6 correct terms. Beyond these two methods, the next-best method selected, on average, 12 correct terms. Variable selection for VAE_JAMOTTON got the worst result, on average, selecting only 7.8 correct terms for the model. We can also see that for almost all methods, variable selection added additional incorrect terms, where only VAE_JAMOTTON picked on average 0.4 additional incorrect terms. These combined facts indicate that using VAE_JAMOTTON results in a small/lean model, which is not desirable in this case.
From the data-specific results, we observe rankings that are almost identical to those for the linear case. This is expected, since we did not introduce any additional inter-variable relationships beyond the interaction between the response and the two covariates. Here again, the use of autoencoders improves MAE/MAPE for categorical columns; however, this improvement is not sufficient to beat the results of MICE-based methods. Again, MICE_TABULATOR shows very strong performance on categorical columns. Additionally, MICE_VV and MICE_TABULATOR indicate a loss of the original data correlation structure.
Upon analyzing synthetic data on a more granular level, we noticed that MICE-based methods can generate numeric values that ignore business constraints, e.g., generating observations with a driver age below $18$. Imposing such constraints may be investigated in a follow-up paper.
Complementing the results in Table 1 and Table 2, we analyzed the averaged (across all experiments) covariance matrices for numeric variables in the generated data. The resulting covariance matrices support the findings above, identifying MICE_PART_SYN and MICE_PART_SYN methods give, on average, the closest covariance matrices to the covariance matrix for the original numerical variables. However, it is important to note that GAN-based models performed better than MICE-based methods with respect to the variance of the density column. In particular, CTGAN maintained accurate variance, whereas the MICE-based methods underestimated this dispersion measure compared to the original data.
@@
Having discussed data generation fidelity, we now focus on data augmentation. For this, we examine how the value of the $M_{1}$ metric depends on the proportion of synthetic data added. We observe that as we increase the proportion of additional synthetic data compared to the original training data, the value for the $M_{1}$ metric increases. The only case in which we observed a decrease in a metric is for the metric $M_{1}$ for the method MICE_PART_SYN and a synthetic data proportion of $0.2$ in the interaction case . In this case, the augmented data achieved the $M_{1}$ metric value of $0.78717$ while the corresponding training data metric was $0.86047$. This is the only result where this happened, it is not enough to say that augmentation of the original freMTPL2freq data might help when it comes to better accuracy of coefficients.
The general relationship between the increase in the proportion of additional synthetic data and the metric $M_{1}$ value is shown in Figure 2. As the proportion of synthetic data increases, we observe a clear upward trend in the $M_{1}$ metric value for the given method. Based on the way the $M_{1}$ metric is calculated and how GLM relies on the underlying data to compute the coefficient values, we estimate that if we keep on increasing the proportion of additional synthetic data, then we will asymptotically reach the metric values which we can compute for the case where we use only synthetic data. These pure synthetic data results are shown on the right side of the plot.